Food Science:
How much low fat milk (1% FAT) and and how much how much whole milk (4% FAT)
should be up to make 5 gal. of reduced FAT milk (2% FAT).
Thank you in advance.

Answer by nerdybill(7003)

How much low fat milk (1% FAT) and and how much how much whole milk (4% FAT)
should be up to make 5 gal. of reduced FAT milk (2% FAT).
.
Let x = amount (gal) of low fat milk
then
5-x = amount (gal) of whole milk
.
.01x + .04(5-x) = .02(5)
.01x + .20-.04x = .10
.20-.03x = .10
-.03x = -.10
x = 3.33 gallons (low fat)
.
whole milk:
5-x = 5-3.33 = 1.67 gallons